River's chemistry class is doing an experiment in which the students create a 30% hydrochloric acid mixture by mixing a certain amount of 25% solution with a different amount of 50% solution. The amount of 25% solution used is 20 liters less than 5 times the amount of 50% solution used. How much of the 25% solution amd the 50% solution are needed in all?
Answer by Boreal(15235)

x=0.25 solution (liters)
y=0.50 solution (liters)
Sense that there should be more of the 25%, because the final solution is 30%, which is closer to 25% than it is to 50%.
5x-20=y
0.25x + 0.50y= 0.30(x+y);;
0.25 x + 0.50 (5x-20)=0.30x+0.30 y
0.25x +2.5x-10=0.30x + 0.30 y;;; 2.75 x -10=0.30x + 1.5x -6. Collect terms
0.95x=4
x=4.21 lit. ANSWER
5x-20= 1.052 l ANSWER

Check
0.25 *4.21 + .5*1.052 should equal 5.262*.3
1.0525+0.526=1.579
5.262*.3=1.1579